# references.bib LLMs left, right, and center: Assessing GPT’s capabilities to label political bias from web domains

### v1

@@ Abstract
This research investigates whether OpenAI’s GPT-4, a state-of-the-art large language model, can accurately classify the political bias of news sources based solely on their URLs. Given the subjective nature of political labels, third-party bias ratings like those from Ad Fontes Media, AllSides, and Media Bias/Fact Check (MBFC) are often used in research to analyze news source diversity. This study aims to determine if GPT-4 can replicate these human ratings on a seven-degree scale ("far-left" to "far-right"). The analysis compares GPT-4’s classifications against MBFC’s, and controls for website popularity using Open PageRank scores. Findings reveal a high correlation (, , ) between GPT-4’s and MBFC’s ratings, indicating the model’s potential reliability. However, GPT-4 abstained from classifying approximately of the dataset, particularly less popular and less biased sources. The study also identifies a slight leftward skew in GPT-4’s classifications compared to MBFC’s. The analysis suggests that while GPT-4 can be a scalable, cost-effective tool for political bias classification of news websites, but its use should complement human judgment to mitigate biases. Further research is recommended to explore the model’s performance across different settings, languages, and additional datasets.
Keywords gpt-4 large language models (LLM) media bias data labeling political bias
1 Introduction
Is The New York Times more politically inclined to the left or the right? The answer might vary according to the respondent. However, it seems to converge to the center-left —at least according to three different services that specialize in rating the news: Ad Fontes Media, AllSides, and Media Bias/Fact Check (MBFC) [AllSidesMediaBias2024, InteractiveMediaBias2023, NewYorkTimes2024]. Finding that sort of information about one of the most-read news outlets in the world is relatively easy, especially given that its public editor acknowledges the newspaper is perceived as left-wing [spaydWhyReadersSee2016]. It might be harder, however, to assess the political bias of more obscure or niche sources where accessible and accurate information is lacking.
While political labels carry a certain degree of subjectivity, third-party bias ratings may help balance information ecosystems by allowing a better understanding of what is consumed, which is particularly relevant since news outlets do not generally declare their standings [sheridanShouldYouTrust2021]. This labeling is often used in academic research. For example, [trielliSearchNewsCurator2019, 8-9] used it to assess whether news content highlighted by Google Search has political leanings. [falkenbergGrowingPolarizationClimate2022] used it to spot an increased polarization on Twitter in the discussions about the United Nations Conference of the Parties on Climate Change, tracking relevant topics for pro-climate users and for those who oppose climate action.
@@ 1 Introduction
Artificial Intelligence (AI) might be an alternative for the labeling process, and recent studies have shown high accuracy when employing Large Language Models (LLMs) for data annotation, including within political contexts [chiangCanLargeLanguage2023, gilardiChatGPTOutperformsCrowd2023, tornbergChatGPT4OutperformsExperts2023, wuLargeLanguageModels2023]. While some recent studies use such models to label news content based on web domain (to assess credibility [yangLargeLanguageModels2023a], for example), none explored judging political bias based on URLs.
Considering this scenario, this research investigates whether OpenAI’s GPT-4, the state-of-the-art language model according to relevant benchmarks [HolisticEvaluationLanguage2024, MMLUBenchmarkMultitask2024], can reliably classify news sources based solely on web domain on a seven-degree scale ("far-left," "left," "center-left," "no bias," "center-right," "right," and "far-right"). It follows the questions:
In addressing the first question, this paper analyzes the correlation between political ideology ratings attributed by GPT and MBFC. The findings reveal high alignment between these ratings, indicating that LLMs may be reliable for labeling political biases based on website domains. The second question relates to the functioning of these models, which involves discerning patterns in extensive text corpora [rooseHowDoesChatGPT2023]. The hypothesis is that how often a news source appears in this data might affect the model’s performance. The study indeed identifies a link: GPT-4 demonstrates a higher likelihood of abstaining from rating less frequently visited websites, which also suffer from less accurate assessments.
The analysis further indicates a tendency of the LLM to avoid classifying sources that MBFC considers to be centrist, resulting in more polarized outputs. It also notes a slight leftward skew in GPT’s classifications compared to MBFC’s. Despite outputting scores that correlate to those assigned by humans, GPT’s classifications are surrounded by a great deal of uncertainty as the criteria they use are unknown, which might hide additional pitfalls. Therefore, while this paper presents encouraging evidence that LLMs can assess the political slant of new sources based on their URLs, users should proceed cautiously, as biases can affect the output. A mixed approach with AI and human ratings is suggested to counter that.
@@ 2.1 GPT-4 and LLMs
[…]
These sorts of AI are created after analyzing the patterns in immense amounts of textual data. They work by predicting combinations of tokens333Word or part of a word the model processes and generates. [omiyeLargeLanguageModels2023, rooseHowDoesChatGPT2023] and reply to questions made by users conversationally, besides following instructions based on their input (called a prompt). LLMs can perform a myriad of language-processing tasks, such as answering questions, writing, summarizing texts, and translating [maslejArtificialIntelligenceIndex2023, 99-108].
As these models grow in terms of computation, number of parameters, and training dataset size, they show capabilities that go beyond their training scope and perform well in tasks that are not their primary goal. These are known as emergent properties. They are not entirely understood but seem related to model size [weiEmergentAbilitiesLarge2022, 2-4]. The lack of clarity about them is an issue, as little is known about their limitations [tornbergChatGPT4OutperformsExperts2023, 4].
@@ 2.2 LLMs for Annotation Tasks
Among LLMs’ emergent properties are applications related to annotating datasets, which motivated the usage investigated in this paper. LLMs were succesfully used to assess the quality of texts written by humans and robots regarding grammar, cohesiveness, likability, and relevance [chiangCanLargeLanguage2023]. Research also shows that GPT in a zero-shot setting can classify tweets according to several criteria, such as whether they are a content moderation issue or relate to pre-defined political topics, with higher accuracy than non-expert humans under the same instructions [gilardiChatGPTOutperformsCrowd2023].
In a similar context to this paper, research found a strong correlation (, ) between human and GPT-assigned ratings for news outlets’ credibility in a corpus of over 7,000 web domains [yangLargeLanguageModels2023a]. The result is based on an aggregate score of multiple services that evaluate news domain quality rather than a single classifying service [linHighLevelCorrespondence2023]. These services correlate with each other on varying degrees ( range: .32-.90), so GPT would fall in the middle of that distribution. This motivated investigating whether GPT could repeat the same level of performance in a political bias classification scenario.
@@ 2.2.1 LLMs in Political Analysis
LLMs have also displayed promising results in tasks related specifically to political analysis. GPT could accurately classify the political affiliation of a Twitter user based on the content of a single anonymized tweet [tornbergChatGPT4OutperformsExperts2023]. The analysis used posts made by all US senators, providing an unequivocal basis for evaluating the answers. The AI performed better than both expert and non-expert humans.
[…]
@@ 2.3 Rating News Sources’ Political Bias
[…]
MBFC, the baseline in this paper, primarily analyzes news source quality. Its database includes information beyond credibility, such as the source’s country and media type (newspaper or TV, for example). It rates political bias by examining editorial content by their position on general philosophy, abortion, economic policy, education policy, environmental policy, gay rights, gun rights, health care, immigration, military, personal responsibility, regulation, social views, taxes, voter ID laws, and workers’ rights [LeftVsRight2021]. It is cited as a source for asserting news quality both in Academia and in the news [akerPredictingNewsSource2019, 2][]NoEvidenceDisease2021[4]resnickIffyQuotientPlatform2018.
@@ 2.3.1 Applications of these Ratings
Analyzing the political bias of news sources is helpful to gauge diversity in the news ecosystem. By classifying ideology and relating it to zip codes, it is possible to note that political ideology in newspapers is more related to appealing to their audience than ownership diversity, as outlets sharing an owner have diverging political stances depending on their readership. This is relevant for regulating the industry (deciding whether to limit ownership, for example) [gentzkowWhatDrivesMedia2010]. These can also be used to measure diversity in news exposure, when, for example, news sources’ political inclinations to analyze the results offered by Google’s algorithm. The classification allowed researchers to identify it was slightly skewed to the left, which might direct readers’ attention that way [trielliSearchNewsCurator2019].
@@ 3 Methodology
The analysis gathered a dataset with classifications from MBFC and popularity ratings from Open PageRank. First, the data for every report made by MBFC was downloaded, removing those that did not include information on political bias, such as the ones focusing on scientific content.
The web domains from MBFC’s database were cleaned to remove inconsistencies (invalid URLs, for example), and missing values were fetched manually from the reports. Two pairs of records shared the same domain, and only one of each was kept. Three records were related to Facebook pages and were removed.
The dataset includes 5,877 observations extracted from MBFC. The bias column from GPT and MBFC was coded into a scale ranging from -3 to +3, meaning "far-left" to "far-right." This makes it possible to calculate both the difference and absolute difference between the two measurements (). See Appendix A for summary statistics.
@@
The analysis primarily focuses on the correlation between the ratings assigned by GPT and MBFC, an approach similar to the work done by [yangLargeLanguageModels2023a] when comparing the LLM’s ability to judge news outlet credibility. There are multiple ways of classifying news sources’ political bias (Section 2.3), and the exact placement in a scale might vary (as there is no clear line separating center-left from left, for example). Thus, this work is concerned with whether GPT will place a source more to the right when MBFC does it, instead of checking whether it predicts the exact value, so correlation is a better measurement. The system’s accuracy is also analyzed later, through linear and logistic regressions controlling for website popularity and political stance.
@@ 3.1 Open PageRank
Open PageRank is a free initiative that applies metrics like the ones used by Google to rank websites, enabling comparison among domains through a scale from 0 to 10 to the decimal level where higher can be interpreted as a proxy for more popular. It provides data about the top 10 million websites in its database [WhatOpenPageRank2023]. The root domain of each URL ("www.example.com/abc" "example.com") is used to retrieve ratings from Open PageRank’s API. Therefore, this analysis does not differentiate between sources that share a root URL ("theguardian.com/observer" and "theguardian.com," for example), meaning popularity might be slightly exaggerated for some. These, however, are not frequent in the data, which contains 5,684 unique root URLs. The distribution of Open PageRank scores is available in Appendix B.
@@ 3.2 Prompting GPT-4
The GPT version used is gpt-4-1106-preview, the most recent available, queried through OpenAI’s API. It was released on Nov. 06, 2023, with up-to-date information until April 2023, and is described by OpenAI as having "improved instruction following" capabilities [NewModelsDeveloper2023]. Being in "preview" means it does not support high traffic [OpenAIPlatform2024], which is not the case in this work. It was cheaper than older versions of GPT-4 (US$0.01/1k input and US$0.03/1k output tokens vs. US$0.03/1k input and US$0.06/1k output tokens) [Pricing2024]. It can reply in JSON format instead of text, making parsing easier as it removes the need to extract the output from a paragraph.
@@
The prompts asked the AI to rate the political bias of a URL or return "-1" if it could not classify it. After the main experiment, to better understand the machine’s behavior, additional prompts for a random sample asked GPT to justify the output. Sample prompts and responses are available in Appendix C. It was used in a zero-shot setting in an attempt to evoke latent information in the model’s training. The request specified that the response be in JSON and provided an example of formatting. A system message added context telling GPT to act as an assistant to determine the political bias of websites.
@@ 4.1 Correlation Between GPT-4 and Human-Assigned Political Bias Ratings
[…]
@@
The correlation between GPT-4 and MBFC’s classifications is at least strong () across most statistically relevant categories. A full breakdown of the categories is available in Appendix D. In terms of country, it was stronger in ones that speak English and weakest in unknown regions (, ) and others (, ). The correlation was statistically relevant in all locations ().
@@
Correlation drops when crunching the different political biases into left, right, and center (including center-left and center-right). It is negligible when considering only sources originally classified as left (, , ) and remains strong with both right (, , ) and center-leaning sources (, , ). That pattern is unexpected, and shows a potential limitation to applying this in practice. It might be due to the loss of nuance when collapsing the scale for this analysis, or due to MBFC’s dataset having more right-wing sources. It could also be due to systematic issues with GPT’s training data. In that case, a possible explanation is that right-wing media is more often directly referred to as that in GPT’s training dataset. This matter requires further investigation, which is beyond the scope of this paper.

### v2

@@ Abstract
This research investigates whether OpenAI’s GPT-4, a state-of-the-art large language model, can accurately classify the political bias of news sources based solely on their URLs. Given the subjective nature of political labels, third-party bias ratings like those from Ad Fontes Media, AllSides, and Media Bias/Fact Check (MBFC) are often used in research to analyze news source diversity. This study aims to determine if GPT-4 can replicate these human ratings on a seven-degree scale ("far-left" to "far-right"). The analysis compares GPT-4’s classifications against MBFC’s, and controls for website popularity using Open PageRank scores. Findings reveal a high correlation (, , ) between GPT-4’s and MBFC’s ratings, indicating the model’s potential reliability. However, GPT-4 abstained from classifying approximately of the dataset. It is more likely to abstain from rating unpopular websites, which also suffer from less accurate assessments. The LLM tends to avoid classifying sources that MBFC considers to be centrist, resulting in more polarized outputs. Finally, this analysis shows a slight leftward skew in GPT’s classifications compared to MBFC’s. Therefore, while this paper suggests that while GPT-4 can be a scalable, cost-effective tool for political bias classification of news websites, its use should be as a complement to human judgment to mitigate biases.
Keywords gpt-4, large language models (LLM), media bias, data labeling, political bias
Is The New York Times more politically inclined to the left or the right? The answer might vary according to the respondent. However, it seems to converge to the center-left —at least according to three different services that specialize in rating the news: Ad Fontes Media, AllSides, and Media Bias/Fact Check (MBFC) [AllSidesMediaBias2024], [InteractiveMediaBias2023], [NewYorkTimes2024]. Finding that sort of information about one of the most-read news outlets in the world is relatively easy, especially given that its public editor acknowledges the newspaper is perceived as left-wing [spaydWhyReadersSee2016]. It might be harder, however, to assess the political bias of more obscure or niche sources where accessible and accurate information is lacking.
While political labels carry a certain degree of subjectivity, third-party bias ratings may help balance information ecosystems by allowing a better understanding of what is consumed, which is particularly relevant since news outlets do not generally declare their standings [sheridanShouldYouTrust2021]. This labeling is often used in academic research. Examples include assessing whether news content highlighted by Google Search has political leanings [trielliSearchNewsCurator2019a, 8-9] and spotting an increased polarization on Twitter in the discussions about the United Nations Conference of the Parties on Climate Change [falkenbergGrowingPolarizationClimate2022].
@@
Artificial Intelligence (AI) might be an alternative for the labeling process, and recent studies have shown high accuracy when employing Large Language Models (LLMs) for data annotation, including within political contexts [chiangCanLargeLanguage2023], [gilardiChatGPTOutperformsCrowd2023], [tornbergChatGPT4OutperformsExperts2023a], [wuLargeLanguageModels2023]. While some recent studies use such models to label news content based on web domain (to assess credibility [yangLargeLanguageModels2023], for example), none explored judging political bias based on URLs.
Considering this scenario, this research investigates whether OpenAI’s GPT-4, the state-of-the-art language model according to relevant benchmarks [HolisticEvaluationLanguage2024], [MMLUBenchmarkMultitask2024], can reliably classify news sources based solely on web domain on a seven-degree scale ("far-left," "left," "center-left," "no bias," "center-right," "right," and "far-right"). It follows the questions:
RQ1
Can an LLM, using only its embedded knowledge with no internet access, classify the political bias of news sources based solely on their URLs in alignment with humans?
RQ2
Does classification accuracy vary depending on the popularity of the news source?
In addressing the first question, this paper analyzes the correlation between political ideology ratings attributed by GPT and MBFC. The second question relates to the functioning of these models, which involves discerning patterns in extensive text corpora [rooseHowDoesChatGPT2023]. The hypothesis is that how often a news source appears in this data might affect the model’s performance.
@@ 1.1 GPT-4 and LLMs
GPT-4 is the latest of OpenAI’s family of LLMs [GPT42023]. It was chosen due to GPT’s popularity in the AI community since the launch of ChatGPT in November 2022 [huChatGPTSetsRecord2023]. It is the best option available111Google’s Gemini Ultra had a higher score but was not publicly available at the time of writing. for the tasks analyzed in this research according to MMLU [MMLUBenchmarkMultitask2024], a widely used benchmark for models in zero-shot settings [hendrycksMeasuringMassiveMultitask2021], [maslejArtificialIntelligenceIndex2023], and HELM, a comprehensive benchmark for language tasks [HolisticEvaluationLanguage2024], [liangHolisticEvaluationLanguage2023], [maslejArtificialIntelligenceIndex2023, 114].
These sorts of AI are created after analyzing the patterns in immense amounts of textual data. They work by predicting combinations of tokens and reply to questions made by users conversationally, besides following instructions based on their input (called a prompt) [omiyeLargeLanguageModels2023], [rooseHowDoesChatGPT2023]. LLMs can perform a myriad of language-processing tasks, such as answering questions, writing, summarizing texts, and translating [maslejArtificialIntelligenceIndex2023, 99-108].
As these models grow in terms of computation, number of parameters, and training dataset size, they show capabilities that go beyond their training scope and perform well in tasks that are not their primary goal. These are known as emergent properties. They are not entirely understood but seem related to model size [weiEmergentAbilitiesLarge2022, 2-4]. The lack of clarity about them is an issue, as little is known about their limitations [tornbergChatGPT4OutperformsExperts2023a, 4].
@@ 1.2 LLMs for Annotation Tasks
Among LLMs’ emergent properties are applications related to annotating datasets. LLMs were successfully used to assess the quality of texts written by humans and robots regarding grammar, cohesiveness, likability, and relevance [chiangCanLargeLanguage2023]. Research also shows that GPT in a zero-shot setting can classify tweets according to several criteria, such as whether they are a content moderation issue or relate to pre-defined political topics, with higher accuracy than non-expert humans under the same instructions [gilardiChatGPTOutperformsCrowd2023].
In a similar context to this paper, research found a strong correlation (, ) between human and GPT-assigned ratings for news outlets’ credibility in a corpus of over 7,000 web domains [yangLargeLanguageModels2023]. The result is based on an aggregate score of multiple services that evaluate news domain quality rather than a single classifying service [linHighLevelCorrespondence2023]. These services correlate with each other on varying degrees ( range: .32-.90), so GPT would fall in the middle of that distribution. This motivated the investigation of whether GPT could repeat the same level of performance in a political bias classification scenario.
@@ 1.2.1 LLMs in Political Analysis
LLMs have also displayed promising results in tasks related specifically to political analysis. GPT could accurately classify the political affiliation of a Twitter user based on the content of a single anonymized tweet [tornbergChatGPT4OutperformsExperts2023a]. The analysis used posts made by all US senators, providing an unequivocal basis for evaluating the answers. The AI performed better than both expert and non-expert humans.
GPT has also shown remarkable capacity to classify US senators in terms of their liberal-conservative ideology, support of gun control, and support of abortion rights, based only on their names and parties using pairwise comparisons [wuLargeLanguageModels2023, 2]. The rankings are not simply mimicking other scales, contrasting with the idea that these systems might be parroting patterns from their training data [benderDangersStochasticParrots2021]. If they were merely copying information from elsewhere, it would be simpler to use the original data. Instead, the investigation shows that the ratings came from a mix of senators’ behaviors and how these politicians are perceived [wuLargeLanguageModels2023, 19]. This indicates that LLMs’ capabilities in political classification tasks warrant further investigation, as they might offer not only a cheaper and faster way of labeling the data but also a new scale altogether.
@@ 1.3 Rating News Sources’ Political Bias
Perception and behavior are currently used forms of rating political bias in media sources. Scales from the likes of AllSides and polling institutes, such as Pew Research Center, classify news sources based on how consumers perceive them [HowAllSidesRates2016], [amymitchellPoliticalPolarizationMedia2014]. Language (behavior) has been used in research to classify newspapers: phrases connected to the right meant right-leaning tendencies [gentzkowWhatDrivesMedia2010]. The measurement correlated to the political incline of readers, given the preferences in the zip codes where the newspapers circulated [gentzkowWhatDrivesMedia2010, 64]. A study used the self-reported ideology of Facebook users to classify news sources by leveraging popularity among the right or the left to gauge the outlet’s political ideology [bakshyExposureIdeologicallyDiverse2015a].
MBFC, the baseline in this paper, primarily analyzes news source quality. Its database includes information beyond credibility, such as the source’s country and media type (newspaper or TV, for example). It rates political bias by examining editorial content by their position on general philosophy, abortion, economic policy, education policy, environmental policy, gay rights, gun rights, health care, immigration, military, personal responsibility, regulation, social views, taxes, voter ID laws, and workers’ rights [LeftVsRight2021]. It is cited as a source for asserting news quality both in Academia and in the news [akerPredictingNewsSource2019, 2], [NoEvidenceDisease2021], [resnickIffyQuotientPlatform2018, 4].
@@ 1.3.1 Applications of these Ratings
Analyzing the political bias of news sources is helpful in gauging diversity in the news ecosystem. By classifying ideology and relating it to zip codes, it is possible to note that political ideology in newspapers is more related to appealing to their audience than ownership diversity, as outlets sharing an owner have diverging political stances depending on their readership. This is relevant for regulating the industry (deciding whether to limit ownership, for example) [gentzkowWhatDrivesMedia2010]. These can also be used to measure diversity in news exposure, when, for example, news sources’ political inclinations to analyze the results offered by Google’s algorithm. The classification allowed researchers to identify it was slightly skewed to the left, which might direct readers’ attention that way [trielliSearchNewsCurator2019a].
@@ 2 Methodology
The analysis gathered a dataset with classifications from MBFC and popularity ratings from Open PageRank. First, the data for every report made by MBFC was downloaded, removing those that did not include information on political bias, such as the ones focusing on scientific content. The web domains from MBFC’s database were cleaned to remove inconsistencies (invalid URLs, for example), and missing values were fetched manually from the reports. Two pairs of records shared the same domain, and only one of each was kept. Three records were related to Facebook pages and were removed.
The dataset includes 5,877 observations extracted from MBFC. The bias column from GPT and MBFC was coded into a scale ranging from -3 to +3, meaning "far-left," "left," "center-left," "center," "center-right," "right," and "far-right." This makes it possible to calculate both the difference and absolute difference between the two measurements (). See the appendices for summary statistics.
@@
The analysis primarily focuses on the correlation between the ratings assigned by GPT and MBFC, an approach similar to previous work that compared the LLM’s ability to judge news outlet credibility [yangLargeLanguageModels2023]. There are multiple ways of classifying news sources’ political bias, and the exact placement in a scale might vary (as there is no clear line separating center-left from left, for example). Thus, this work is concerned with whether GPT will place a source more to the right when MBFC does it instead of checking whether it predicts the exact value, so the correlation is a better measurement. The system’s accuracy is also analyzed later through linear and logistic regressions controlling for website popularity and political stance.
@@ 2.1 Open PageRank
Open PageRank is a free initiative that applies metrics like the ones used by Google to rank websites, enabling comparison among domains through a scale from 0 to 10 to the decimal level where higher can be interpreted as a proxy for more popular. It provides data about the top 10 million websites in its database [WhatOpenPageRank2023]. The root domain of each URL ("www.example.com/abc" "example.com") is used to retrieve ratings from Open PageRank’s API. Therefore, this analysis does not differentiate between sources that share a root URL ("theguardian.com/observer" and "theguardian.com," for example), meaning popularity might be slightly exaggerated for some. These, however, are not frequent in the data, which contains 5,684 unique root URLs. The distribution of Open PageRank scores is available in the appendices.
@@ 2.2 Prompting GPT-4
The GPT version used is gpt-4-1106-preview, the most recent available when writing, queried through OpenAI’s API. It was released on Nov. 06, 2023, with up-to-date information until April 2023, and is described by OpenAI as having "improved instruction following" capabilities [NewModelsDeveloper2023]. Being in "preview" means it does not support high traffic [OpenAIPlatform2024], which is not the case in this work. It was cheaper than older versions of GPT-4 (US$0.01/1k input and US$0.03/1k output tokens vs. US$0.03/1k input and US$0.06/1k output tokens) [Pricing2024]. It can reply in JSON format instead of text, making parsing easier as it removes the need to extract the output from a paragraph.
@@
The prompts asked the AI to rate the political bias of a URL or return "-1" if it could not classify it. After the main experiment, to better understand the machine’s behavior, additional prompts for a random sample asked GPT to justify the output. Sample prompts and responses are available in the appendices. It was used in a zero-shot setting in an attempt to evoke latent information in the model’s training. The request specified that the response be in JSON and provided an example of formatting. A system message added context telling GPT to act as an assistant to determine the political bias of websites.
@@ 3.1 Correlation Between GPT-4 and Human-Assigned Political Bias Ratings
The analysis shows a very strong222Based on Quinnipiac University’s table for interpreting Spearman’s correlation coefficients in Politics contexts [akogluUserGuideCorrelation2018, 92] correlation between GPT-4’s and MBFC’s political bias classifications (, , ). This shows that both vary in the same direction: sources classified toward one side of the spectrum in MBFC are likely to follow the same pattern with GPT.
@@
The correlation between GPT-4 and MBFC’s classifications is at least strong () across most statistically relevant categories. A full breakdown of the categories is available in the appendices. In terms of country, it was stronger in ones that speak English and weakest in unknown regions (, ) and others (, ). The correlation was statistically relevant in all locations ().
@@
Correlation drops when crunching the different political biases into left, right, and center (including center-left and center-right). It is negligible when considering only sources originally classified as left (, , ) and remains strong with both right (, , ) and center-leaning sources (, , ). That pattern is unexpected, and shows a potential limitation to applying this in practice. It might be due to the loss of nuance when collapsing the scale for this analysis or due to MBFC’s dataset having more right-wing sources. It could also be due to systematic issues with GPT’s training data. In that case, a possible explanation is that right-wing media is more often directly referred to as that in GPT’s training dataset. This matter requires further investigation, which is beyond the scope of this paper.